Is the matrix
A =


1 1 1
2 3 0
3 8 3

 .
invertible? (5)
4.2 Solve the following system by Gauss-Jordan elimination
x + y + z = 2
2x + 3y + 2z = 5
3x + 8y + z = 11
(10)
4.3 Now solve the system in 4.2 by applying Cramer’s Rule.
1
Expert's answer
2020-04-23T18:50:19-0400

It is necessary to determine whether the matrix

"A=\begin{pmatrix}
 1 & 1&1 \\
 2 & 3&0 \\
 3 & 8& 3 
\end{pmatrix}"

is invertible.

To calculate the inverse matrix, we write down matrix A with additional identity matrix:

"\left(\begin{matrix}
1 & 1 & 1 \vert& 1 & 0 & 0 \\
2 & 3 & 0\vert & 0 & 1 & 0 \\
3 & 8 & 3\vert & 0 & 0 & 1
\end{matrix}\right)"

To find the inverse matrix, we will use elementary transformations over the rows of the matrix, turn the left part of the resulting matrix into a single one.

Subtract the 1st row, multiplied by 2; subtract the 1st row multiplied by 3 from the 3rd row

"\left(\begin{matrix}
1 & 1 & 1& \vert& 1 & 0 & 0 \\
0 & 1& -2&\vert & -2 & 1 & 0 \\
0 &5 & 0&\vert & -3 & 0 & 1
\end{matrix}\right)"

subtract the second row multiplied by 1 from the 1st row; subtract the 2nd row multiplied by 5 from the 3rd row

"\left(\begin{matrix}
1 & 0 & 3& \vert& 3 & -1 & 0 \\
0 & 1& -2&\vert & -2 & 1 & 0 \\
0 &0 & 10&\vert & 7 & -5 & 1
\end{matrix}\right)"

The 3rd line is divisible by 10

"\left(\begin{matrix}
1 & 0 & 3& \vert& 3 & -1 & 0 \\
0 & 1& -2&\vert & -2 & 1 & 0 \\
0 &0 & 1&\vert & 0.7 & -0.5 & 0.1
\end{matrix}\right)"

subtract from the 1st row the 3rd row multiplied by 3; to the 2nd row add the 3rd row multiplied by 2

"\left(\begin{matrix}
1 & 0 & 0& \vert& 0.9 & 0.5 & -0.3 \\
0 & 1& 0&\vert & -0.6 & 0 & 0.2 \\
0 &0 & 1&\vert & 0.7 & -0.5 & 0.1
\end{matrix}\right)"

Answer:

"\left(\begin{matrix}
1 & 0 & 0& \vert& 0.9 & 0.5 & -0.3 \\
0 & 1& 0&\vert & -0.6 & 0 & 0.2 \\
0 &0 & 1&\vert & 0.7 & -0.5 & 0.1
\end{matrix}\right)"

"A^{-1}=\begin{pmatrix}
 0.9 & 0.5&-0.3 \\
 -0.6 &0 &0.2 \\
 0.7 & -0.5& 0.1 
\end{pmatrix}"

4.2 Solve the following system by eliminating Gauss-Jordan

x + y + z = 2

2x + 3y + 2z = 5

3x + 8y + z = 11 

Expansion of the matrix of systems to the step form is given:


"\begin{pmatrix}
 1 & 1&1&\vert2 \\
 2 & 3&2 &\vert5 \\
 3 & 8& 1&\vert11 
\end{pmatrix}
R_2-2*R_1\mapsto R_2
\begin{pmatrix}
 1 & 1&1&\vert2 \\
 0 & 1&0 &\vert1 \\
 3 & 8& 1&\vert11 
\end{pmatrix}
R_3-3*R_1\mapsto R_3\\
\begin{pmatrix}
 1 & 1&1&\vert2 \\
 0 & 1&0 &\vert1 \\
 0 &5& -2&\vert5 
\end{pmatrix}
R_3-5*R_2\mapsto R_3 
\begin{pmatrix}
 1 & 1&1&\vert2 \\
 0 & 1&0 &\vert1 \\
 0 & 0& -2&\vert0 
\end{pmatrix}
\frac{R_3}{-2}\mapsto R_3\\
\begin{pmatrix}
 1 & 1&1&\vert2 \\
 0 & 1&0 &\vert1 \\
 0 & 0& 1&\vert0 
\end{pmatrix}
R_1-1*R_3\mapsto R_1\\ 
\begin{pmatrix}
 1 & 0&0&\vert1 \\
 0 & 1&0 &\vert1 \\
 0 & 0& 1&\vert0 
\end{pmatrix}"

"\begin{cases}
 x=1 \\
 y=1 (1)\\
z=0
\end{cases}"


From equation 3 of system (1) we find the variable z

z=0

From equation 2 of system (1) we find the variable y

y=1

From equation 1 of system (1) we find the variable x

x=1

Answer: x=1; y=1; z=0.

The overall solution "X=\begin{pmatrix}
 1 \\
 1 \\
 0 \\
\end{pmatrix}"

4.3 Now solve the system in 4.2, applying the Cramer's rule.


"\begin{cases}
 x+y+z=2 \\
 2x+3y+2z=5 \\
3x+8y+z=11
\end{cases}\\
\Delta=\begin{vmatrix}
 1 & 1&1 \\
 2 & 3&2\\
3&8&1
\end{vmatrix}=-2 \\
\Delta_1=\begin{vmatrix}
 2 & 1&1 \\
 5 & 3&2\\
11&8&1
\end{vmatrix}=-2 \\

\Delta_2=\begin{vmatrix}
 1 & 2&1 \\
 2 & 5&2\\
3&11&1
\end{vmatrix}=-2 \\
\Delta_3=\begin{vmatrix}
 1 & 1&2 \\
 2 & 3&5\\
3&8&11
\end{vmatrix}=0 \\
x=\frac{\Delta_1}{\Delta}=\frac{-2}{-2}=1\\
y=\frac{\Delta_2}{\Delta}=\frac{-2}{-2}=1\\
z=\frac{\Delta_3}{\Delta}=\frac{0}{-2}=0\\"


Answer: x=1; y=1; z=0